Solve the system of linear equations using the Graphing Method.
2x+y=5
3x-2y=4

Answer by haileytucki(390)

2x+y=5_3x-2y=4
Since 2x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2x from both sides.
y=-2x+5_3x-2y=4
Since 3x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3x from both sides.
y=-2x+5_-2y=-3x+4
Divide each term in the equation by -2.
y=-2x+5_-(2y)/(-2)=-(3x)/(-2)+(4)/(-2)
Simplify the left-hand side of the equation by canceling the common factors.
y=-2x+5_y=-(3x)/(-2)+(4)/(-2)
Simplify the right-hand side of the equation by simplifying each term.
y=-2x+5_y=(3x)/(2)-2
Create a graph to locate the intersection of the equations. The intersection of the system of equations is the solution.
y=-2x+5_y=(3x)/(2)-2